What is the enthalpy change if 105 g of ethane is burned in a combustion reaction whose molar enthalpy of combustion is -1251 kJ/mol. *


1
Expert's answer
2021-04-29T07:36:24-0400

n = m / M

M (C2H6) = 30.07 g/mol

n (C2H6) = 105 / 30.07 = 3.5 mol

delta H (C2H6) = - 1251 x 3.5 = -4367 kJ